solve for x (3x+2)^2+(3x-2)^2=26
Answer by unlockmath(1688)

Hello,
Let's expand these out:
(3x+2)^2+(3x-2)^2=26
9x^2+12x+4+9x^2-12x+4=26
Combine like terms:
18x^2+8=26
Subtract 8 and divide by 18:
x^2=18/18 or 1
Sq rt both sides:
x=+-1
Make sense?
RJ
www.math-unlock.com